# Solving the Bongard-logo Problem by Modeling a Probabilistic Model

### v7

@@ Abstract
Abstract reasoning problems pose challenges to the perception and cognition abilities of AI algorithms, demanding deeper pattern recognition and inductive reasoning beyond mere identification of explicit image features. In this study, we introduce PMoC, a probabilistic model tailored for the Bongard-Logo problem, achieving high reasoning accuracy through the construction of an conditional probabilistic model. Additionally, we have designed the Pose-Transformer, an enhanced Transformer-Encoder specifically crafted for complex abstract reasoning tasks, including Bongard-Logo, RAVEN, I-RAVEN, and PGM. Inspired by the pose matrix in capsule networks, Pose-Transformer strengthens the focus on positional relationships between local features when processing image data. When combined with PMoC, it can further enhance reasoning accuracy. Our Pose-Transformer effectively addresses reasoning difficulties associated with changes in the position of abstract entities, outperforming previous models on RAVEN’s OIG, D3x3 subsets, and the PGM dataset. Finally, considering the deployment difficulties arising from the large number of Pose-Transformer parameters, this paper presents a lightweight version, Straw-Pose-Transformer, which maintains performance while significantly reducing the parameter count. This study contributes to enhancing AI capabilities in abstract reasoning and cognitive pattern recognition.
@@
Abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Probabilistic Model.
@@ I Introduction
Deep neural networks have achieved remarkable success in various fields, including computer vision [1, 3, 2], natural language processing [4, 5, 6], generative models [8, 7, 9], visual question answering [10, 11], and abstract reasoning[15, 16, 17]. Deep learning, as a pivotal branch of machine learning, Simulate the learning process of the human brain by establishing multilayered neural networks, thereby facilitating the learning and inference of intricate patterns within data. In the domain of graphical abstract reasoning, deep learning has found extensive applications to address diverse and intricate pattern recognition and reasoning problems[15, 16, 17].
Through deep learning, we can architect multilayered neural network models that, by undergoing extensive training and learning from vast datasets, can autonomously learn and extract pertinent features and patterns. These extracted features and patterns enable operations such as classification, regression, and clustering on novel data, facilitating automated decision-making and inference. The applications of deep learning are vast and widespread, positioning it as a pivotal research direction within the artificial intelligence sphere. However, research and applications of deep learning are not devoid of challenges and issues, encompassing aspects like model generalizability, training stability, and efficiency.
Graphical reasoning problems, in particular, pose multidimensional challenges to deep learning. The challenges that graphical reasoning problems present to deep learning are multifaceted and often distinct from conventional machine learning tasks.
Firstly, graphical reasoning problems demand that models should capture and decipher inherent regularities and patterns within graphs. However, due to the intricacy and variability of graphs, models must possess highly flexible and scalable capabilities to adapt to diverse graphical structures and characteristics. This aspect poses a significant challenge to existing deep learning models due to their often-limited architectural rigidity and representational capacity.
Secondly, graphical reasoning problems require models that can handle the compositionality and structure of graphs. This necessitates the model’s capacity to comprehend the constituent elements of graphs and their interrelationships for effective reasoning and prediction. However, prevalent deep learning models often face impediments in addressing such problems due to their limited ability to explicitly model and efficiently handle graphical structures.
Additionally, graphical reasoning problems mandate models to possess abstraction and generalization capabilities for graphs. This entails the model’s proficiency in extracting general regularities and patterns from specific graphical instances for subsequent reasoning and judgment on novel graphs. Nevertheless, deep learning models, which typically undergo training on specific instances, often exhibit subpar performance in abstraction and generalization tasks.
Furthermore, addressing graphical reasoning problems necessitates models with efficient learning and reasoning capabilities. Given the complexity and large-scale nature of graphical data, models must be adept at rapid learning and reasoning within constrained computational resources for real-time applicability and scalability in practical scenarios. However, prevalent deep learning models often encounter challenges like high computational complexity and prolonged training durations when dealing with large-scale graphical data, posing significant challenges in tackling graphical reasoning problems.
Consequently, addressing the challenges posed by graphical reasoning problems to deep learning constitutes a pivotal research direction. This necessitates efforts towards designing more flexible and scalable deep learning models, enhancing representational and learning methodologies for models, and developing efficient learning and reasoning algorithms. By surmounting these challenges, we can propel the application and advancement of deep learning in graphical reasoning problems, providing smarter and more efficient solutions for a myriad of practical problems.
@@
For instance, problems like Ravens Progressive Matrices (RPM)[15] and Bongard problems[16, 17] present learning requirements that span from perception to reasoning.
I-A The RAVEN database
As exemplified in figure1, the construction of RAVEN problems adheres to a structured framework. Within this framework, human-defined concepts inherent to geometric images, such as ”shape” and ”color”, undergo a process of artificial abstraction. This abstraction transforms these concepts into bounded, countable, and precise ”visual attribute” values. Subsequently, the notion of ”rule” is invoked to delineate the systematic progression of a circumscribed set of ”visual attribute” values. It is noteworthy that certain visual attributes remain unconstrained by the ”rule”, potentially introducing elements of interference in the reasoning processes of deep models.
Figure 1: RAVEN case
The generative process for a complete RAVEN problem commences with the selection of a rule sample from a predefined rule repository. This is followed by the design of visual attribute values aligned with the contents of the chosen ”rule”. Attributes not governed by the rule undergo random value assignments. Subsequently, image rendering occurs based on the generated ”attribute” information.
The RAVEN database exhibits diversity through its multiple sub-databases. These include single-rule groups: center single (center), distribute four (G2$\times$2), distribute nine (G3$\times$3), and double-rule groups: in center single out center single (O-IC), up center single down center single (U-D), left center single right center single (L-R), in distribute four out center single (O-IG). In problems adhering to single-rule variations, the progressive attribute changes within image entities are bound by a unitary set of rules. Conversely, double-rule variations enforce constraints via two independent sets of rules, augmenting the complexity of the problem.
I-B Pgm database
Pgm[19] and RAVEN problems exhibit a striking parallelism in their design philosophy, with both frameworks employing a problem statement consisting of 8 images and an answer pool supplemented by an additional 8 images. A unique facet of PGM problems is the expansion of the ”rule” concept, which not only encompasses the progressive patterning of ”visual attributes” horizontally within the matrix but also vertically. An illustrative example of a PGM problem is presented in the accompanying figure 2, highlighting its intricate nature.
Figure 2: PGM case
The intricacy of RPM problems, therefore, arises not only from the extraction of visual attributes across multiple hierarchies but also from the challenge of inducing and learning the progressive patterns displayed by these ”visual attributes.” This emphasizes the significance of discerning not only the attributes themselves but also their structural progression within these problems, a task that requires a nuanced understanding of both the individual components and their interrelationships within the problem context.
I-C Bongard-logo database
@@ II Related work
II-A Graphical abstract reasoning dataset
The original RAVEN dataset[18] possessed a significant flaw in its candidate answer generation process. Specifically, each incorrect answer within the candidate set had only one erroneous visual attribute value, with the rest being identical to the correct answer. Consequently, the most frequent value among the candidate answers for each visual attribute would correspond to the correct visual attribute value. As a result, models could achieve remarkably high accuracy by merely considering the answer pool as input, without accounting for the question matrix’s statement section. This rendered the performance on the RAVEN database an inadequate reflection of the model’s true capabilities. To address this limitation, I-RAVEN[20] and RAVEN Fair[21] were developed as remedies.
In the realm of computer vision, the success of deep learning models often hinges on extensive data availability. However, the initial Bongard problems comprised only 100 instances, and even subsequent expansions left the question count relatively limited. Given that deep learning’s remarkable strides in computer vision are largely data-dependent, Nie et al. developed a procedurally generated technique to address the data scarcity issue, resulting in the Bongard-Logo[16] benchmark database for Bongard problems. This comprehensive database encompasses 2000 Bongard problems, categorized into three main types: free-form problems, basic-shape problems, and abstract-shape problems.
In free-form problems, the shape patterns within the test set exhibit longer strokes compared to those in the training set. The basic-shape problem test set poses challenges in recognizing combinations of two shapes, with these foundational shapes appearing only once in the training set, ensuring no conceptual overlap between the test and training sets. The abstract-shape problem test set is further bifurcated: the combined abstract shape test set and the novel abstract shape test set [12, 13, 14].
Within the combined abstract shape test set, 20 pairwise combinations of two abstract attributes are randomly selected, with 20 problems per combination. While all individual attributes in this test set have been observed separately in the training set, these 20 novel combinations are exclusive to the test set. This necessitates the model’s ability to comprehend abstract shape concepts and utilize them in combinations. On the other hand, the novel abstract shape test set extracts one attribute from the training set along with all its combinations with other attributes. All questions related to the retained attribute are unique in this test set. The specialized test sets demand that models not merely memorize shape concepts but also demonstrate combinatorial generalization and extrapolation capabilities [12, 13, 14]. These rigorous test sets aim to evaluate the model’s profound understanding of shape concepts and its innovative prowess.
II-B RPM solver
@@
The CoPINet model[22], for instance, introduces a novel contrast module that enables the network to learn subtle differences between input graphs. This is complemented by an inference module that distills potential underlying rules from the graph structures. By contrast, the LEN+teacher model[23] leverages a student-teacher paradigm to determine optimal training sequences and make predictions based on this guided learning.
The DCNet model[24] takes a different approach, employing a dual contrast module to directly compare rule rows and columns. This allows the model to pinpoint differences among candidate answers more precisely. Meanwhile, the NCD model[25] eschews traditional supervised learning methods, instead opting for an unsupervised approach that incorporates pseudo-targets and decentralization techniques to train the network.
In the context of subgraph-level analysis, SCL[26] establishes multiple monitors for subgraphs within reasoning problems. This ensures that each branch of the model can focus on specific visual attributes or rules, enhancing the model’s ability to handle complex reasoning tasks. SAVIR-T [27] further extends this concept by extracting information not only within subgraphs but also between them, providing a more holistic view of the reasoning problem and thus improving reasoning capabilities.
Recent research has also highlighted the benefits of using relatively decoupled perceptual visual features for reasoning tasks. These features, when combined with symbolic attention methods, offer not only higher reasoning accuracy but also stronger model interpretability. This is exemplified in the PrAE[28] model, where neurosymbolic systems incorporate prior knowledge of rules to perform probabilistic reasoning and generate answer images.
On the other end of the spectrum, models like ALANS[29] eliminate the need for prior rule knowledge entirely. Despite this, they demonstrate superior generalization capabilities compared to end-to-end models. This underscores the importance of balancing learning and inference in complex reasoning tasks.
NVSA[30], employs holographic vectorized representations along with ground-truth attribute values to construct a neural-symbolic model, aiming to enhance the model’s comprehension and reasoning capabilities. This approach combines the strengths of both neural networks and symbolic logic, leveraging their respective advantages in handling continuous data patterns and discrete symbolic rules.
Finally, RS-CNN[31] and RS-TRAN[31] combine the best aspects of the aforementioned methods to achieve state-of-the-art results on RPM problems. RS-TRAN, in particular, leverages the powerful Transformer-Encoder architecture to tackle RPM problems from multiple perspectives and through multiple inferences. Meanwhile, RS-CNN adopts a multi-scale and multi-inference approach that combines both local and global information for more accurate reasoning[37]. In this paper, we have designed a novel standard layer that can replace the conventional Transformer-Encoder within RS-Tran. When this new layer is integrated, RS-Tran exhibits significant improvement in inference progress, demonstrating its enhanced efficiency and performance.
Overall, these models represent the cutting edge of image reasoning research and offer promising directions for future work.
II-C Bongard solver
@@
II-D Capsule Network
Capsule Networks (CapsNets)[33] are a novel deep neural network architecture. The design inspiration for this model comes from the “capsule” structures in the human brain, which are able to effectively process different types of visual stimuli and encode information such as position, shape, and speed.
In the context of deep learning, a capsule is a collection of embedded neurons, and a CapsNet is composed of capsules rather than individual neurons. It operates through a dynamic routing algorithm to update network weights, enabling the network to better understand and recognize features in images. Compared to traditional Convolutional Neural Networks (CNNs), CapsNets demonstrate stronger feature representation capabilities and improved performance.
Specifically, each capsule in a CapsNet represents a specific feature detector. The output of a capsule is a vector, where the length of the vector represents the probability of the feature’s existence, and the direction of the vector encodes the pose information of the feature. As a feature moves within the network, the output vector of the capsule changes accordingly, while maintaining its length (i.e., the probability of the feature’s existence).
Furthermore, CapsNets utilize an agreement-based dynamic routing algorithm as an alternative to the Max-Pooling operation in traditional CNNs. This dynamic routing algorithm can automatically adjust the structure and connection weights of the network based on the characteristics of the current input data, thereby enhancing the network’s performance and generalization capabilities.
The Capsule Network learns corresponding pose matrices for all local representations of an image, which are utilized to map these local representations onto high-dimensional representations within the object space. Through the Gaussian Routing algorithm, the Capsule Network computes the weights of these mapped high-dimensional representations. These weights are then applied to calculate a weighted sum of the high-dimensional representations, with the compressed summation result indicating pose consistency among local features within the image. During the training process, the pose matrices adapt to learn the relative positional relationships among local representations.
Currently, CapsNets are primarily applied in the field of image recognition and have demonstrated impressive performance on several benchmark tests.
II-E Transformer and Vision Transformer (ViT)
@@ II-F Sinkhron distance
The Sinkhorn algorithm, closely associated with this distance, iteratively updates two sets of scaling factors until convergence to a solution that minimizes the regularized transport cost. This approach is particularly advantageous in high-dimensional settings or when dealing with large datasets, as it avoids the expensive linear programming computations required by the standard Wasserstein distance. The Sinkhorn distance offers a means to measure the similarity between two distributions solely based on samples drawn from them, without the need to have explicit knowledge of their underlying forms. This advantage sets it apart as a powerful tool in scenarios where direct access to the distributional forms is not feasible.
@@ III Methodology for Bongard-Logo
The Bongard-logo[16] and Raven’s Progressive Matrices (RPM)[15] tasks are both paradigms of abstract reasoning that necessitate the extractor’s adeptness in unveiling precise concepts veiled within abstract representations. These concepts are indicative of more elevated abstractions, surpassing rudimentary ideas such as “pixel configuration patterns”. Frequently, these elevated abstractions encapsulate diverse degrees of human-centered presupposed knowledge, including elements like form, dimension, hue, spatial orientations, concave and convex contours of shapes, and shape completion. Through the employment of distinct querying approaches, the Bongard-logo and RPM tasks endeavor to evaluate the extractor’s expertise in unveiling and assimilating these abstract concepts.
This paper attempts to propose a methodology that relies on analyzing the distribution of images in the primary and auxiliary groups of Bongard-Logo problems to solve them. This methodology aims to provide a new direction different from the previous ones for solving Bongard-Logo problems.
@@ III-A A method for solving the bongard-logo problem based on sinkhorn distance (SBSD)
Directly deriving the image distributions $p_{i}(x|y)$ and $q_{i}(x|y)$ of Bongard-Logo without dimensionality reduction and computing the cross measure between these two distributions is quite challenging. Moreover, such an approach would also result in a waste of resources and time. Therefore, we aim to devise a deep learning algorithm, denoted as $f_{\theta}(z|x)$, which can map samples $x_{ij}$ to latent variables $z_{ij}$. We hope that $f_{\theta}(z|x)$ can serve as a distribution transformation function, and by measuring the distribution of the latent representation, we can address the Bongard-Logo problem. Accordingly, the key objective is to ensure that the encoder $f_{\theta}(z|x)$ possesses a specific image encoding capability, which entails producing encoded latent variables that demonstrate specific characteristics. Specifically, the distribution of latent variables within the primary group (represented as $p_{i}^{\prime}(z|y)$) and the distribution of latent variables within the auxiliary group (represented as $q_{i}^{\prime}(z|y)$) should exhibit very low cross-measure. The Kullback-Leibler (KL) divergence, a measure used to quantify the similarity between distributions, can be employed for this purpose[7]. The formula for calculating KL divergence is provided below.
@@
Hence, we endeavored to utilize baseline models, specifically ResNet18, for the purpose of encoding images within the Bongard-Logo paradigm, denoted as $\{x_{ij}\}_{j=1}^{14}$, into meaningful representations, designated as $\{z_{ij}\}_{j=1}^{14}$. These representations, in the context of a Bongard-Logo question, were interpreted as samples drawn from the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$. Subsequently, constraints were imposed on the Sinkhorn distance between $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ through these representations. This methodology was designed primarily because the Sinkhorn algorithm computes the distance without requiring explicit knowledge of the distributions’ specific forms, making it highly suitable for scenarios where the forms of $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ remain unknown.
@@
In summary, SBSD treats the representations of images within the primary group as samples from the primary group distribution $p_{i}^{\prime}(z|y)$, and the auxiliary group follows the same approach. By utilizing these samples, the Sinkhorn distance is employed to calculate and represent the distance between the primary $p_{i}^{\prime}(z|y)$ and auxiliary distributions $q_{i}^{\prime}(z|y)$. This distance can be used to solve the Bongard-Logo problem.

### v8

# Solving the Clustering Reasoning Problems by Modeling a Deep-Learning-Based Probabilistic Model

@@ Abstract
Visual abstract reasoning problems pose significant challenges to the perception and cognition abilities of artificial intelligence algorithms, demanding deeper pattern recognition and inductive reasoning beyond mere identification of explicit image features. Research advancements in this field often provide insights and technical support for other similar domains. In this study, we introduce PMoC, a deep-learning-based probabilistic model, achieving high reasoning accuracy in the Bongard-Logo, which stands as one of the most challenging clustering reasoning tasks. PMoC is a novel approach for constructing probabilistic models based on deep learning, which is distinctly different from previous techniques. PMoC revitalizes the probabilistic approach, which has been relatively weak in visual abstract reasoning. As a bonus, we also designed Pose-Transformer for complex visual abstract reasoning tasks. Inspired by capsule networks, it focuses on positional relationships in image data, boosting accuracy when combined with PMoC. Our Pose-Transformer effectively addresses reasoning difficulties associated with changes in the position of entities, outperforming previous models on RAVEN dataset, and the PGM dataset. RAVEN and PGM represent two significant progressive pattern reasoning problems. Finally, considering the deployment difficulties of Pose-Transformer, we introduced Straw-Pose-Transformer, a lightweight version. This study contributes to enhancing the capabilities of artificial intelligence in abstract reasoning, cognitive pattern, and probabilistic modeling of complex systems.
@@
Visual abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Deep-Learning-Based Probabilistic Model.
@@ I Introduction
Deep neural networks have achieved remarkable success in various fields, including computer vision [1, 3, 2], natural language processing [4, 5, 6], generative models [8, 7, 9], visual question answering [10, 13], and abstract reasoning[14, 15, 16]. Deep learning, as a pivotal branch of machine learning, simulate the learning process of the human brain by establishing multilayered neural networks, thereby facilitating the learning and inference of intricate patterns within data[11]. Deep learning has found extensive applications to address diverse and intricate pattern recognition and reasoning problems[14, 15, 16].
Visual abstract reasoning problems occupy a pivotal position in the field of artificial intelligent. They involve identifying and comprehending hidden patterns, structures, and relationships within graphics, reflecting a deep cognitive process. Research on these problems has propelled advancements in artificial intelligent technology, fostering continuous optimization of key techniques such as feature extraction and modeling. The outcomes of this research have broad applications, spanning security, healthcare, and autonomous driving, among others. Furthermore, the capability of effective reasoning is a significant indicator of the intelligence level of AI systems. Addressing these problems poses challenges but also presents tremendous opportunities for breakthroughs, especially with the advent of deep learning techniques[12]. In summary, visual abstract reasoning problems are not only crucial to basic research but also hold the potential to drive technological progress and expand the application scope of artificial intelligent.
@@
Visual abstract reasoning problems typically come in two forms: clustering reasoning and progressive pattern reasoning. For instance, Ravens Progressive Matrices (RPM)[14] serves as a benchmark for progressive pattern reasoning problems, while Bongard problems[15, 16] stand as a paragon for clustering inference, both encompassing learning requirements that span from perception to reasoning.
I-A Bongard-logo database
@@
I-B The RAVEN database
As exemplified in figure2, the construction of RAVEN problems adheres to a structured framework. Within this framework, human-defined concepts[11] inherent to geometric images, such as “shape” and “color”, undergo a process of artificial abstraction. This abstraction transforms these concepts into bounded, countable, and precise “visual attribute” values. Subsequently, the notion of “rule” is invoked to delineate the systematic progression of a circumscribed set of “visual attribute” values. It is noteworthy that certain visual attributes remain unconstrained by the “rule”, potentially introducing elements of interference in the reasoning processes of deep models.
Figure 2: RAVEN case
The RAVEN database showcases diversity with various sub-databases, including single-rule groups like center single, distribute four (G2x2), distribute nine (G3x3), and double-rule groups such as O-IC, U-D, L-R, and O-IG. Single-rule variations involve progressive attribute changes bound by one rule set, while double-rule variations add complexity with two independent rule sets.
I-C Pgm database
Pgm[18] and RAVEN problems exhibit a striking parallelism in their design philosophy, with both frameworks employing a problem statement consisting of 8 images and an answer pool supplemented by an additional 8 images. A unique facet of PGM problems is the expansion of the “rule” concept, which not only encompasses the progressive patterning of “visual attributes” horizontally within the matrix but also vertically. An illustrative example of a PGM problem is presented in the accompanying figure 3, highlighting its intricate nature.
Figure 3: PGM case
@@ II Related work
II-A RPM solver
@@
The CoPINet model[21] uses a contrast module to learn graph differences and an inference module to extract rules. LEN+teacher[22] employs a student-teacher approach for guided learning. DCNet[23] compares rule rows and columns with a dual contrast module. NCD[24] adopts unsupervised learning with pseudo-targets.
For subgraph analysis, SCL[25] sets up monitors to focus on specific visual attributes or rules. SAVIR-T[26] considers both within and between subgraph information. Recent studies show the benefits of decoupled visual features and symbolic attention for reasoning, as seen in PrAE[27].
ALANS[28] demonstrates strong generalization without prior rule knowledge. NVSA[29] uses holographic representations and ground-truth attributes for neural-symbolic reasoning.
RS-CNN[30] and RS-TRAN[30] combine various methods for RPM problems. RS-TRAN uses a Transformer-Encoder for multi-perspective inference, while RS-CNN integrates local and global information[41]. This paper introduces a new layer for RS-Tran, enhancing its efficiency and performance.
In summary, these models advance image reasoning research and pave the way for future progress.
II-B Bongard solver
@@
II-C Capsule Network
Capsule Networks (CapsNets)[32] are a new type of deep neural network inspired by the “capsule” structures in the human brain, which process visual information efficiently. In CapsNets, neurons are grouped into capsules, and these networks use a dynamic routing algorithm to update weights, improving image feature recognition. Compared to traditional CNNs, CapsNets offer stronger feature representation and better performance. Each capsule detects a specific feature, outputting a vector where length indicates feature existence probability and direction encodes feature pose. CapsNets replace CNNs’ Max-Pooling with an agreement-based dynamic routing, adjusting network structure and weights based on input data, boosting performance and generalization. They also learn pose matrices for image representations, using a Gaussian Routing algorithm to weight high-dimensional representations, indicating pose consistency. CapsNets are mainly used in image recognition and have shown strong performance in tests.
II-D Transformer and Vision Transformer (ViT)
@@ II-E Sinkhron distance
II-F Discrete-Time Inference and Establishment of Contextual Probability Models
The theory of isochronous constraint reasoning [36] has brought profound insights to the design of reasoning problem solvers. By integrating logical reasoning from this theory, along with optimization strategies and discrete-time system analysis, we can more precisely understand and predict complex logical relationships in reasoning problem. Additionally, drawing on the experience of isochronous constraint reasoning in addressing nonlinear problems, and adopting data-driven techniques, we can build an efficient and accurate reasoning problem solver[37], effectively tackling challenges in visual entities and elements sequence reasoning.
The contextual analysis[38] method provides significant insights into the design of abstract reasoning problem solvers. By deeply analyzing the contextual relationships between graphics, the solver can more accurately grasp the logical connections and patterns of change among them. Specifically, this method urges consideration of the position and role of graphics in the overall sequence, leading to a more comprehensive understanding of each graphic’s significance. Encouraging a focus on the mutual influence among graphics rather than viewing each in isolation, the solver’s design benefits from capturing and utilizing contextual information, thereby enhancing problem-solving accuracy and efficiency, especially when dealing with complex and nonlinear graphic relationships. In conclusion, the contextual analysis approach serves as a powerful thinking tool for the design of visual abstract reasoning problem solvers, facilitating the development of smarter and more efficient algorithms.
Attention-based visual models[39] demonstrate significant superiority by focusing on critical information within images, thereby enhancing processing efficiency and accuracy. These models enable a sharper focus on pertinent data while adapting to multi-modal data relationships, resulting in their widespread application across various domains. It also serves as the underpinning of our methodological approach.
In light of the challenges and opportunities associated with modeling system problem spaces using probabilistic models[40], this paper opts to leverage deep learning algorithms to construct a deep-learning-based probabilistic model aimed at addressing visual abstract reasoning problems.
@@ III Methodology for Bongard-Logo
The Bongard-logo[15] and Raven’s Progressive Matrices (RPM)[14] tasks are both paradigms of abstract reasoning that require the extractor’s ability in unveiling precise concepts hidden within abstract representations. These concepts are indicative of more elevated abstractions, surpassing rudimentary ideas such as “pixel configuration patterns”. Frequently, these elevated abstractions encapsulate diverse degrees of human-centered pre-existing knowledge knowledge, including elements like form, dimension, hue, spatial orientations, concave and convex contours of shapes, and shape completion. Using distinct querying approaches, the Bongard-logo and RPM tasks endeavor to evaluate the extractor’s expertise in unveiling and assimilating these abstract concepts.
The methodology presented in this section targets clustering reasoning problems within visual abstract reasoning, attempting to solve them by modeling a probabilistic model. The Bongard-Logo problem stands as a renowned example of clustering reasoning, drawing significant attention due to its exceptional reasoning difficulty and the current lack of progress in solving it. This paper aims to achieve promising results on the Bongard-Logo problem, demonstrating the potential of probabilistic models in solving clustering reasoning tasks.
@@ III-A A method for solving the bongard-logo problem based on sinkhorn distance (SBSD)
Directly deriving the image distributions $p_{i}(x|y)$ and $q_{i}(x|y)$ of Bongard-Logo without dimensionality reduction and computing the cross measure between these two distributions is quite challenging. Moreover, such an approach would also result in a waste of resources and time. Therefore, we aim to devise a deep learning algorithm, denoted as $f_{\theta}(z|x)$, which can map samples $x_{ij}$ to latent variables $z_{ij}$. We hope that $f_{\theta}(z|x)$ can serve as a distribution transformation function, and by measuring the distribution of the latent representation, we can address the Bongard-Logo problem. Accordingly, the key objective is to ensure that the encoder $f_{\theta}(z|x)$ possesses a specific image encoding capability, which necessitates its ability to produce encoded latent variables that demonstrate specific characteristics. Specifically, the distribution of latent variables within the primary group (represented as $p_{i}^{\prime}(z|y)$) and the distribution of latent variables within the auxiliary group (represented as $q_{i}^{\prime}(z|y)$) should exhibit very low cross-measure. The Kullback-Leibler (KL) divergence, a measure used to quantify the similarity between distributions, can be employed as an optimization objective for $f_{\theta}(z|x)$[7]. The formula for calculating KL divergence is provided below.
@@
Hence, we endeavored to utilize baseline models, specifically ResNet18, for the purpose of encoding images within the Bongard-Logo paradigm, denoted as $\{x_{ij}\}_{j=1}^{14}$, into meaningful representations, designated as $\{z_{ij}\}_{j=1}^{14}$. These representations, in the context of a Bongard-Logo question, can be interpreted as samples drawn from the distributions $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$. Subsequently, constraints were imposed on the Sinkhorn distance between $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ through these representations. This methodology was designed primarily because the Sinkhorn algorithm computes the distance without requiring explicit knowledge of the distributions’ specific forms, making it highly suitable for scenarios where the forms of $p_{i}^{\prime}(z|y)$ and $q_{i}^{\prime}(z|y)$ remain unknown.
@@
In summary, SBSD represents our initial and most direct attempt to model the Bongard-Logo problem as a probabilistic framework. SBSD treats the representations of images within the primary group as samples from the primary group distribution $p_{i}^{\prime}(z|y)$, and the auxiliary group follows the same approach. By utilizing these samples, the Sinkhorn distance is employed to calculate and represent the distance between the primary $p_{i}^{\prime}(z|y)$ and auxiliary distributions $q_{i}^{\prime}(z|y)$. This distance can be used to solve the Bongard-Logo problem.